0.6 -2y + 14.4 = 9
Solve...I really need to know how to solve this..

Found 2 solutions by ReadingBoosters, stanbon:
Answer by ReadingBoosters(3246)

Group like terms
-2y + .6 + 14.4 = 9
-2y + 15 = 9
-2y + 15 - 15 = 9 - 15
-2y = -6
-2y/-2 = -6/-2
y = 3

Answer by stanbon(75887)

0.6 -2y + 14.4 = 9
-----
To get rid of the decimals multiply thru by 10:
6 - 20y + 144 = 90
------
-20y + 150 = 90
---
-20y = -60
y = 3
================
Cheers,
Stan H.